What mass of Al2O3 is produced when 13.1 L of O2 gas react with Al at 482 K and 3.40 atm. 

4Al(s)+3O2(g)⟶2Al2O3(g)

 

First, use the ideal gas equation to calculate the moles of O2. 

Then, calculate the mass of Al2O3 from the moles of O2. Show the conversions required to solve this problem.


1
Expert's answer
2022-04-14T04:59:04-0400

pV = nRT

R = 0.082 L atm K−1 mol−1

n = pV / RT = (3.40 x 13.1) / (0.082 x 482) = 1.1 mol

According to the equation, n (Al2O3) = 2/3 x n (O2) =2/3 x 1.1 = 0.75 mol

n = m / M

M (Al2O3) = 101.96 g/mol

m (Al2O3) = n x M = 0.75 x 101.96 = 76.6 g